OpenStudy (anonymous):

rationalize the denominator: 14 / cube root(4x^2)

OpenStudy (anonymous):

\[14*\sqrt[3]{4x^2}*\sqrt[3]{4x^2}\div \sqrt[3]{4x^2}*\sqrt[3]{4x^2}*\sqrt[3]{4x^2}\] \[14*\sqrt[3]{4x^2}*\sqrt[3]{4x^2} \div 4x^2\] \[14*\sqrt[3]{16x^3 x} \div 4x^2\] \[14*\sqrt[3]{16x} \div 4x^2\]